OpenStudy (anonymous):

1. Use the fi rst principles formula lim h -0 f(x + h) f(x) h to find the derivative function for the following function: f(x) = -x^2 + 2x^3

OpenStudy (anonymous):

\[f'(x)=\lim_{h\to0}\frac{\left(2(x+h)^3-(x+h)^2\right)-\left(2x^3-x^2\right)}{h}\] Expand those binomials. Most terms will disappear, leaving you with a factor of \(h\) in both numerator and denominator.

OpenStudy (anonymous):

check the attachment and tell me if this is correct

OpenStudy (anonymous):

Yes, that's all correct.

OpenStudy (anonymous):

thank you @SithsAndGiggles

OpenStudy (anonymous):

You're welcome